OpenStudy (anonymous):

FOR A MEDAL!!!!!!!!!!!! A loan of 44,000 is made at 6.25% interest, compounded annually. After how many years will the amount due reach 63,000

OpenStudy (anonymous):

\[63 = 44(1 + 0.0625)^t\]solve for t

OpenStudy (anonymous):

\[63=44(1.0625)^t\]Solve for t.

OpenStudy (anonymous):

I guess Euler types faster than I do.

OpenStudy (anonymous):

how would i even solve for t? @AnimalAin @Euler271

OpenStudy (anonymous):

I should think the solution might go something like this:\[63=44(1.0625)^t \implies \frac{63}{44}=1.0625^t \implies \ln \frac{63}{44}=\ln1.0625^t\]\[\implies \ln \frac{63}{44}=t \ln1.0625 \implies \frac{ \ln \frac{63}{44}}{ \ln1.0625}=t\]From there, you just plug it into a calculator if you need an approximate answer.